OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = Eight divided by x squared. + 4

OpenStudy (anonymous):

f(x)=1/x g(x) = x^2 +4

OpenStudy (anonymous):

g(x) = 8/x +4 f(x) =8/x +4